# A Digital Twin of Evaporative Thermo-Fluidic Process in Fixation Unit of DoD Inkjet Printers

### v1

@@ 3.4 Comparison of the fixation process model with machine data
This section compares simulation results with sensor data for two test cases: unprinted 350 $\mathrm{g/m^{2}}$ paper and printed 115 $\mathrm{g/m^{2}}$ paper. Unprinted paper involves only linear heat diffusion, while printed paper includes moisture and non-linear evaporation effects, providing comprehensive coverage of the fixation process. Two experimental data sets are used: one for parameter tuning to minimize root mean squared error (RMSE) against the mean sensor data, and the other for validation against simulation results. The mean of the first sensor reading is used as the initial condition for both.
3.4.1 Simulation results for $\mathrm{g/m^{2}}$ blank paper:
@@ 4 Digital Twining of Fixation Process through $\mathcal{H}_{\infty}$ Optimal State Estimator Synthesis
    $\begin{bmatrix}-\gamma I&-\mathscr{D}_{11}^{T}&-(\mathscr{P}\mathscr{B}_{1}+\mathscr{Z}\mathscr{D}_{21})^{T}\\ (\cdot)^{*}&-\gamma I&\mathscr{C}_{1}\\ (\cdot)^{*}&(\mathscr{P}\mathscr{A}+\mathscr{Z}\mathscr{C}_{2})^{T}&(\cdot)^{*}\end{bmatrix}\leq 0.$
@@
The above optimization problem, including the inversion of the operator $\mathscr{P}$, can be solved in PIETOOLS with the help of Linear Matrix Inequalities.
@@ 4.1 $\mathcal{H}_{\infty}$ optimal estimator for blank paper
Initial condition Error signal
peak error

### v2

@@ 3.5 Comparison of the fixation process model with machine data
This section compares simulation results with sensor measurements for two test cases, unprinted 350 $\mathrm{g/m^{2}}$ paper and printed 115 $\mathrm{g/m^{2}}$ paper. The first case involves only linear heat diffusion, while the second additionally includes moisture transport and nonlinear evaporation, providing comprehensive validation of the thermo-fluidic model. Two experimental data sets are used, one for parameter tuning based on RMSE against the mean sensor data, and another for validation. As only the paper’s top surface temperature and the conveyor’s bottom surface are measurable during operation, validation necessarily focuses on these quantities. Real-time moisture measurements are not available in practice, making temperature the only ground truth information for assessing the accuracy of the digital twin. The reliability of the reconstructed moisture profile, therefore, depends on accurately estimating the temperature distribution that drives the evaporation dynamics via the nonlinear maps $\Delta_{1}$ and $\Delta_{2}$. In this sense, the estimator acts as a virtual sensor that uses available temperature measurements to infer the internal temperature and moisture fields that cannot be directly observed.
3.5.1 Simulation results for 350 $\mathrm{g/m^{2}}$ blank paper:
@@ 4 Digital Twining of Fixation Process through $\mathcal{H}_{\infty}$ Optimal State Estimator Synthesis
    $\begin{bmatrix}-\gamma I&-\mathscr{D}_{11}^{*}&-(\mathscr{P}\mathscr{B}_{1}+\mathscr{Z}\mathscr{D}_{21})^{*}\\ (\cdot)^{*}&-\gamma I&\mathscr{C}_{1}\\ (\cdot)^{*}&(\mathscr{P}\mathscr{A}+\mathscr{Z}\mathscr{C}_{2})^{*}&(\cdot)^{*}\end{bmatrix}\leq 0.$
@@
The inequalities in Proposition 3 are operator inequalities where all the operators are PI operators and the unknown PI operators ($\mathscr{P},\mathscr{Z}$) appear linearly in the inequalities. In such a case, these operator inequalities (as well as a convex optimization problem subject to such inequalities) can be solved using Linear Matrix Inequalities (LMIs). Formulation of these LMIs, solving them, and inversion of $\mathscr{P}$ are performed in the PIETOOLS software package.
@@ 4.1 $\mathcal{H}_{\infty}$ optimal estimator for blank paper
Initial condition Error signal peak error (C)